Analysis Price on: JNJ MSFT T WFC WMT XOM 1/2/2013 70.84 27.62 35 35.05 69.24 88.71 12/31/2014 104.57 46.45 33.59 54.82 85.88 92.45 Growth 47.61% 68.18% -4.03% 56.41% 24.03% 4.22% In the cells below, determine how much $100 of each stock would be worth on 12/31/2014 if you purchased it on 1/2/2013. Balance
Answer #1

Answer

1/2/2013 data 70.84 27.62 35 35.05 69.24 88.71

12/31/2014 data 104.57 46.45 33.59 54.82 85.88 92.45

Net Growth 47.61% 68.18% -4.03% 56.41% 24.03% 4.22%

$100 of JNJ would be equal to $100 + 47.61% growth = $100 + {($100*47.61)/100} = $100 + $47.61 = $147.61

$100 of MSFT would be equal to $100 + 68.18% growth = $100 + {($100*68.18)/100} = $100 + $68.18 = $168.18

$100 of T would be equal to $100 -4.03% growth = $100 - {($100*4.03)/100} = $100 - $4.03 = $95.97

$100 of WFC would be equal to $100 + 56.41% growth = $100 + {($100*56.41)/100} = $100 + $56.41 = $156.41

$100 of WMT would be equal to $100 + 24.03% growth = $100 + {($100*24.03)/100} = $100 + $24.03 = $124.03

$100 of XOM would be equal to $100 + 4.22% growth = $100 + {($100*4.22)/100} = $100 + $4.22 = $104.22

Stock name $100 worth on 12/31/2014
JNJ $147.61
MSFT $168.18
T $95.97
WFC $156.41
WMT $124.03
XOM $104.22